# ul. Pasteura 5, 02-093 Warsaw, Poland

### v1

@@ 4.2 Frequency dependence: varying tension
$\rho(t)=\begin{cases}\rho_{\Delta}\left(\frac{a_{\Delta}}{a(t)}\right)^{m}\,,\qquad t<t_{\Delta}\\ \rho_{\Delta}\left(\frac{a_{\Delta}}{a(t)}\right)^{n}\,,\qquad t\geq t_{\Delta}\end{cases}\,,\qquad G\mu(t)=\begin{cases}G\mu_{\Delta}\left(\frac{t_{\Delta}}{t}\right)^{p}\,,\qquad t<t_{\Delta}\\ G\mu_{\Delta}\left(\frac{t_{\Delta}}{t}\right)^{q}\,,\qquad t\geq t_{\Delta}\end{cases}\,.$ (77)
Particularly interesting are cases $p=1$, $q=2/3$, $m=6$, $n\in\{3,4\}$ and $p=2/3$, $q=0$, $m\in\{3,4\}$ and $n=3$ since these correspond, respectively, to a kination period followed by a tracker, and the end of a (matter or radiation) tracker followed by a period of matter domination given by oscillations of the modulus responsible for varying the string tension (the volume modulus being a good candidate for these choices of $p$ and $q$).
In this case the results are different due to the evolution equation of the loops (20)-(21), which we re-write here for convenience:
$\ell(t)=\left(\frac{t}{t_{i}}\right)^{p/2}\left(\alpha P^{\chi}-\frac{\Gamma G\mu_{0}}{\frac{3p}{2}-1}\left(\frac{t_{0}}{t_{i}}\right)^{p}\right)t_{i}+\frac{\Gamma G\mu_{0}}{\frac{3p}{2}-1}\left(\frac{t_{0}}{t}\right)^{p-1}t_{0},\qquad(p\neq 2/3),$ (78)
$\ell(t)=\left[\alpha P^{\chi}-\Gamma G\mu_{0}\left(\frac{t_{0}}{t_{i}}\right)^{2/3}\log\left(\frac{t}{t_{i}}\right)\right]\left(\frac{t}{t_{i}}\right)^{1/3}t_{i}\,,\qquad(p=2/3)\,,$ (79)
@@
$\left(\frac{t}{t_{i}}\right)^{1-3p/2}\ll\frac{\alpha P^{\chi}}{\Gamma G\mu(t_{i})}\,.$ (80)
@@
$\frac{\ell(t)}{t_{\Delta}}\simeq\alpha P^{\chi}\left(\frac{t}{t_{\Delta}}\right)^{p/2}\left(\frac{t_{i}}{t_{\Delta}}\right)^{1-p/2}\,\rightarrow\frac{t_{i}}{t_{\Delta}}\simeq\left(\frac{f}{f_{\Delta}}\right)^{2/(p-2)}\left(\frac{t}{t_{\Delta}}\right)^{(4/m-p)/(2-p)}\,,$ (81)
$\frac{dt_{i}}{d\ell(t)}\simeq\frac{2}{(2-p)\alpha P^{\chi}}\left(\frac{t_{i}}{t}\right)^{p/2}\,.$ (82)
@@
$\frac{\ell(t)}{t_{\Delta}}\simeq\alpha P^{\chi}\left(\frac{t}{t_{\Delta}}\right)^{q/2}\left(\frac{t_{i}}{t_{\Delta}}\right)^{1-p/2}\,\rightarrow\frac{t_{i}}{t_{\Delta}}=\left(\frac{f}{f_{\Delta}}\right)^{2/(p-2)}\left(\frac{t}{t_{\Delta}}\right)^{(4/n-q)/(2-p)}\,.$ (83)
$\frac{dt_{i}}{d\ell(t)}\simeq\frac{2}{(2-p)\alpha P^{\chi}}\left(\frac{t_{i}}{t_{\Delta}}\right)^{p/2}\left(\frac{t_{\Delta}}{t}\right)^{q/2}\,.$ (84)
These quantities are all we need to compute the GW spectrum sourced by these loops (noting, however, that depending in the details of the cosmic evolution this spectrum might always be subdominant). Let us begin by computing the contribution at times $t<t_{\Delta}$. Introducing the above expressions into Eq. (63), we find:
$h^{2}\Omega_{\rm{GW}}^{(1)}\simeq\frac{2A_{\Delta}}{2-p}\left(\frac{f}{f_{\Delta}}\right)^{D}\int_{t_{F}}^{t_{\Delta}}{\frac{dt}{t_{\Delta}}\,\left(\frac{t}{t_{\Delta}}\right)^{E}}$ (85)
$D=\frac{6(m-2)}{m(2-p)}\,,\qquad E=\frac{4}{m}-\frac{5p}{2}+\frac{4-p\,m}{m(2-p)}\left(\frac{6}{m}+\frac{p}{2}-4\right)\,,$ (86)
which reproduces the low-frequency results of the section above (where the length-loss due to GW emission is negligible) for $p=0$. The behaviour is now qualitatively different because in all cases of interest $m=6,\,p=1$ and $m\in\{3,4\}\,,p=2/3$ the quantity $E$ is smaller or equal than $-1$ and the spectrum is dominated by emission at early times. The integral is thus dominated at the earliest times, where the loops of size $\ell(t_{\rm{min}}(f))$ are formed. This $t_{\rm{min}}(f)$ can be computed by requiring $t_{i}(t_{\rm{min}}(f),f)=t_{\rm{min}}$, finding [^31] This is true provided $f<f_{F}$. For $f>f_{F}$ the condition $t_{i}(t_{\rm{min}}^{\prime}(f),f)>t_{F}$ is the relevant one, which provides a new $t_{min}(f)^{\prime}$ given by $t_{\rm{min}}(f)^{\prime}/t_{F}\geq(f/f_{F})^{-2m/(4-pm)}\,,$ For $p\,m>4$, there is no spectrum at frequencies larger than $f<f_{F}$ (at least sourced by the first mode). Otherwise the condition can be satisfied and gives the cutoff at higher frequencies, resulting in a different spectral index. We will be interested in $p\,m>4$ in what follows.
$\frac{t_{\rm{min}}}{t_{F}}=\left(\frac{f}{f_{F}}\right)^{-m/(m-2)}\,.$ (87)
@@
$h^{2}\Omega_{\rm{GW}}^{(1)}\simeq A_{\Delta}\left(\frac{f}{f_{\Delta}}\right)^{F}\,,\qquad F=\frac{2(m(p+1)-4)}{m-2}\,.$ (88)
@@
$f_{\Delta}<f<f_{\Delta}\left(\frac{t_{\Delta}}{t_{F}}\right)^{(m-2)/m}=f_{F}\,.$ (89)
@@
$h^{2}\Omega_{\rm{GW}}^{\rm{(1,peak)}}\simeq A_{\Delta}\left(\frac{t_{kin}}{t_{F}}\right)^{2(1+p-4/m)}\,.$ (90)
@@
$D=\frac{6(m-2)}{m(2-p)}\,,\qquad E=\frac{4}{n}-\frac{5q}{2}+\frac{4-q\,n}{n(2-p)}\left(\frac{6}{m}+\frac{p}{2}-4\right)\,,$ (91)
The integral is again dominated by early times (namely $E<-1$) for kination (regardless of whether it is followed by a matter or radiation period, this being a tracker or not), while for the tracker periods it is dominated by late times. Thus, in the case where only the tracker is present one would need to compare the early and late time contribution to compute the frequency behaviour, etc. We postpone this analysis for future work and focus in the case of volume modulus kination for concreteness, where the earliest times dominate. Let us also point out, however, that in these scenarios the spectral index will be generically different than in scenarios with constant tension. In the case of volume modulus kination, for instance, $\Omega_{\rm{GW}}\sim f^{4}$, resulting in a very large boost in the signal.
@@ 5 Benchmark scenarios
We now turn to illustrating the GW spectra arising from cosmic superstrings in the two scenarios described in Sec. 3.1.2. We operate under the assumption that the PTA signal is explained by cosmic superstrings, and search for further consequences of this assumption. Following Ref. Ellis:2023tsl , where $\beta=1/2$ and $\chi=0$, the preferred string tension and intercommutation probability are [^32] We note here that in Avgoustidis:2025svu the choice is $\beta=2/3$ and $\chi\neq 0$, which is why they find a different best-fit than in Ellis:2023tsl .
$G\mu=2\times 10^{-12}\,,\qquad P=4\times 10^{-3}\,.$ (92)
Using Eqs. (1) and (4), this fixes the volume and the string coupling to
$\mathcal{V}=3.3\times 10^{9}\,,\qquad g_{s}=0.11\,,$ (93)
@@
The vanilla LVS makes a concrete prediction in the LISA band in terms of the mass and decay rate of the volume mode. The transition frequency depends on model-dependent order one numbers, with the most likely scenario being that the LISA mission would be able to measure a deviation from standard cosmology. Due to the large suppression of the signal from a long period of early matter domination, there is no meaningful enhancement in the high-frequency portion of the spectrum (despite an epoch of early kination). The SVL-LVS has an interesting behaviour: the high-frequency signal grows for larger hierarchies between the inflation and late-time vacuum, with a long period of kination in between. As we will discuss, it is not easy to explain the PTA signal whilst being consistent with a large signal at high-frequencies. We nevertheless provide an example where such a large, high-frequency signal can occur as an illustration of the interesting physics behind the fact that the signals grow with large hierarchies. The basic physics can be applied to other well-motivated scenarios which can potentially lead to large signals (see for example Brunelli:2025ems ).
@@ 5.1 Vanilla LVS and LISA
Consider the mass of the volume modulus in LVS Conlon:2005ki : [^33] Following Ebelt:2023clh , we define $W_{0}=\sqrt{\frac{2}{\pi}}|e^{K_{\rm{cs}}/2}W_{\rm{GVW}}|$ in terms of the Kähler potential of the complex structure moduli and the Gukov-Vafa-Witten Gukov:1999ya superpotential. This is because Ebelt:2023clh finds a $W_{0}$ following a Gaussian distribution around 0.
$m_{\Phi}=\mathcal{O}(1)\times\frac{W_{0}}{\mathcal{V}^{3/2}g_{s}^{1/4}}M_{P}$ (94)
which results in a gravitational decay rate, using Eq. (41):
$\Gamma_{\Phi}=\frac{\mathcal{O}(1)}{48\pi}\cdot\frac{W_{0}^{3}}{\mathcal{V}^{9/2}g_{s}^{3/4}}M_{P}\,.$ (95)
This decay rate then translates into a time for decay of the modulus, which in turn predicts a deviation from the flat spectrum at a frequency that we compute as follows. Because, as discussed in Sec. 4.1, the largest contribution to the amplitude at a frequency $f$ from loops that are formed during matter but source during radiation arises from the latest times, the spectrum at $f$ is dominated by the loops that are formed and source during radiation domination. Out of these, the largest contribution arises at $t_{M}(f)$ as defined in Eq. (72). The transition from the flat spectrum sourced by loops that are created and emit during radiation domination is therefore predicted to occur when $t_{i}(t_{M}(f),f)>1/\Gamma_{\Phi}$, meaning that no loop created during radiation domination sources at its point of maximum emission $t_{M}(f)$ in this frequency bin. In formulae, using Eqs. (68) and (72), this implies a transition $f_{\rm{T}}$ given by:
$f_{\rm{T}}=f_{\Gamma}\left(\frac{\alpha P^{\chi}}{\Gamma G\mu}\right)^{1/2}=\frac{2}{\left(\alpha P^{\chi}\times\Gamma G\mu\right)^{1/2}}\frac{\Gamma_{\Phi}}{z_{\Gamma}}\,,$ (96)
where the second equality defines $f_{\Gamma}$ as a particular case of $f_{\Delta}$ (cf: Eq. (68)) in terms of the time of reheating $t_{\Gamma}=1/\Gamma_{\Phi}$ and its redshift, $z_{\Gamma}$. Assuming instantaneous reheating and standard cosmology after volume decay:
$f_{\rm{T}}=8.6\times 10^{-4}\text{ Hz }\,\times\,\left(\frac{0.1}{\alpha P^{\chi}}\times\frac{2\times 10^{-12}}{G\mu}\times\frac{c_{i}}{48\pi}\right)^{1/2}\times\left(\frac{m_{\Phi}}{30\text{ TeV}}\right)^{3/2}\,,$ (97)
@@
Figure 2 illustrates the difference between the spectrum for standard cosmology and in vanilla LVS for the benchmark value $m_{\Phi}=30$ TeV, corresponding in LVS to the values in Eq. (93), together with $W_{0}\simeq 1$ and modulo order one numbers. The Figure also illustrates the mass-dependence of the frequency with the mass of the modulus. A lighter volume mode corresponds to a lower-frequency signal (since the decay occurs later in time), although in vanilla scenarios these values might be in tension with achieving a successful Nucleosynthesis, as discussed in Sec. 3.1.2. Ignoring this issue, a lighter volume mode can be achieved by choice of $W_{0}$ or by a stabilization mechanism leading to $m_{\Phi}\sim 1/\mathcal{V}^{a}$ with $a>3/2$. A heavier volume, on the other hand, corresponds to a larger transition frequency. Microscopically, this could be obtained for large values of $W_{0}$ (though note that $W_{0}$ is bounded by above for consistent description of the effective theory, see e.g. Cicoli:2013swa ), or by a stabilization mechanism similar to the one discussed in Appendix A. In addition, Figure 3 zooms into the interferometer band.
Figure 2: GW spectrum sourced by cosmic superstrings with $G\mu=2\times 10^{-12}$ and $P=4\times 10^{-3}$, in the vanilla LVS scenario, discussed in Section 5.1. A period of early matter domination induced by the volume modulus ends by gravitational decay. This induces a dip in the spectrum that will be robustly probed by the LISA mission. A reasonable estimate of the volume mass is 30 TeV (solid black) and is related to the dip frequency by Eq. (97). We also show the spectrum in “Standard” cosmology, i.e: radiation domination with constant tension.
@@
There are a multitude of presently operating and proposed GW experiments that could detect the SGWB studied in this Section [^34] We have used the publicly available data from Fu:2024rsm to plot the experimental sensitivities. (see Caprini:2018mtu for a review in sources and detection efforts of cosmological GW backgrounds and Aggarwal:2020olq ; Aggarwal:2025noe for proposals at frequencies above the kHz). In the plots, we show sensitivity curves for the following detectors:
Figure 3: GW spectrum sourced by cosmic superstrings with $G\mu=2\times 10^{-12}$ and $P=4\times 10^{-3}$ in the large-scale interferometer band for the vanilla LVS scenario, discussed in Section 5.1. A period of early matter domination induced by the volume modulus ends by gravitational decay. This induces a dip in the spectrum that will be robustly probed by the LISA mission. A reasonable estimate of the volume mass is 30 TeV (solid black) and is related to the dip frequency by Eq. (97). We also show the spectrum in “Standard” cosmology, i.e: radiation domination with constant tension.

### v2

@@ 4.2 Frequency dependence: varying tension
$\rho(t)=\begin{cases}\rho_{\Delta}\left(\frac{a_{\Delta}}{a(t)}\right)^{m}\,,\qquad t<t_{\Delta}\\ \rho_{\Delta}\left(\frac{a_{\Delta}}{a(t)}\right)^{n}\,,\qquad t\geq t_{\Delta}\end{cases}\,,\qquad G\mu(t)=\begin{cases}G\mu_{\Delta}\left(\frac{t_{\Delta}}{t}\right)^{p}\,,\qquad t<t_{\Delta}\\ G\mu_{\Delta}\left(\frac{t_{\Delta}}{t}\right)^{q}\,,\qquad t\geq t_{\Delta}\end{cases}\,.$ (4.19)
Particularly interesting are the cases $p=1$, $q=2/3$, $m=6$, $n\in\{3,4\}$ and $p=2/3$, $q=0$, $m\in\{3,4\}$ and $n=3$ since these correspond, respectively, to a kination period followed by a tracker, and the end of a (matter or radiation) tracker followed by a period of matter domination given by oscillations of the modulus responsible for varying the string tension (the volume modulus being a good candidate for these choices of $p$ and $q$).
In this case the results are different due to the evolution equation of the loops (2.20)-(2.21), which we re-write here for convenience:
$\ell(t)=\left(\frac{t}{t_{i}}\right)^{p/2}\left(\alpha P^{\chi}-\frac{\Gamma G\mu_{0}}{\frac{3p}{2}-1}\left(\frac{t_{0}}{t_{i}}\right)^{p}\right)t_{i}+\frac{\Gamma G\mu_{0}}{\frac{3p}{2}-1}\left(\frac{t_{0}}{t}\right)^{p-1}t_{0},\qquad(p\neq 2/3),$ (4.20)
$\ell(t)=\left[\alpha P^{\chi}-\Gamma G\mu_{0}\left(\frac{t_{0}}{t_{i}}\right)^{2/3}\log\left(\frac{t}{t_{i}}\right)\right]\left(\frac{t}{t_{i}}\right)^{1/3}t_{i}\,,\qquad(p=2/3)\,$, (4.21)
@@
$\left(\frac{t}{t_{i}}\right)^{1-3p/2}\ll\frac{\alpha P^{\chi}}{\Gamma G\mu(t_{i})}\,.$ (4.22)
@@
$\frac{\ell(t)}{t_{\Delta}}\simeq\alpha P^{\chi}\left(\frac{t}{t_{\Delta}}\right)^{p/2}\left(\frac{t_{i}}{t_{\Delta}}\right)^{1-p/2}\,\rightarrow\frac{t_{i}}{t_{\Delta}}\simeq\left(\frac{f}{f_{\Delta}}\right)^{2/(p-2)}\left(\frac{t}{t_{\Delta}}\right)^{(4/m-p)/(2-p)}\,$, (4.23)
$\frac{dt_{i}}{d\ell(t)}\simeq\frac{2}{(2-p)\alpha P^{\chi}}\left(\frac{t_{i}}{t}\right)^{p/2}\,.$ (4.24)
@@
$\frac{\ell(t)}{t_{\Delta}}\simeq\alpha P^{\chi}\left(\frac{t}{t_{\Delta}}\right)^{q/2}\left(\frac{t_{i}}{t_{\Delta}}\right)^{1-p/2}\,\rightarrow\frac{t_{i}}{t_{\Delta}}=\left(\frac{f}{f_{\Delta}}\right)^{2/(p-2)}\left(\frac{t}{t_{\Delta}}\right)^{(4/n-q)/(2-p)}\,.$ (4.25)
$\frac{dt_{i}}{d\ell(t)}\simeq\frac{2}{(2-p)\alpha P^{\chi}}\left(\frac{t_{i}}{t_{\Delta}}\right)^{p/2}\left(\frac{t_{\Delta}}{t}\right)^{q/2}\,.$ (4.26)
These quantities are all we need to compute the GW spectrum sourced by these loops (noting, however, that depending in the details of the cosmic evolution this spectrum might always be subdominant). Let us begin by computing the contribution at times $t<t_{\Delta}$. Introducing the above expressions into Eq. (4.5), we find:
$h^{2}\Omega_{\rm{GW}}^{(1)}\simeq\frac{2A_{\Delta}}{2-p}\left(\frac{f}{f_{\Delta}}\right)^{D}\int_{t_{F}}^{t_{\Delta}}{\frac{dt}{t_{\Delta}}\,\left(\frac{t}{t_{\Delta}}\right)^{E}}$ (4.27)
$D=\frac{6(m-2)}{m(2-p)}\,,\qquad E=\frac{4}{m}-\frac{5p}{2}+\frac{4-p\,m}{m(2-p)}\left(\frac{6}{m}+\frac{p}{2}-4\right)\,$, (4.28)
which reproduces the low-frequency results of the section above (where the length-loss due to GW emission is negligible) for $p=0$. The behaviour is now qualitatively different because in all cases of interest m=6,p=1m=6,\,p=1 and $m\in\{3,4\}\,,p=2/3$ the quantity $E$ is smaller or equal than 1-1 and the spectrum is dominated by emission at early times. The integral is thus dominated at the earliest times, where the loops of size $\ell(t_{\rm{min}}(f))$ are formed. This $t_{\rm{min}}(f)$ can be computed by requiring $t_{i}(t_{\rm{min}}(f),f)=t_{\rm{min}}$, finding [^34] This is true provided $f<f_{F}$. For $f>f_{F}$ the condition $t_{i}(t_{\rm{min}}^{\prime}(f),f)>t_{F}$ is the relevant one, which provides a new $t_{min}(f)^{\prime}$ given by $t_{\rm{min}}(f)^{\prime}/t_{F}\geq(f/f_{F})^{-2m/(4-pm)}\,$, For pm>4p\,m>4, there is no spectrum at frequencies larger than $f<f_{F}$ (at least sourced by the first mode). Otherwise the condition can be satisfied and gives the cutoff at higher frequencies, resulting in a different spectral index. We will be interested in pm>4p\,m>4 in what follows.
$\frac{t_{\rm{min}}}{t_{F}}=\left(\frac{f}{f_{F}}\right)^{-m/(m-2)}\,.$ (4.29)
@@
$h^{2}\Omega_{\rm{GW}}^{(1)}\simeq A_{\Delta}\left(\frac{f}{f_{\Delta}}\right)^{F}\,,\qquad F=\frac{2(m(p+1)-4)}{m-2}\,.$ (4.30)
@@
$f_{\Delta}<f<f_{\Delta}\left(\frac{t_{\Delta}}{t_{F}}\right)^{(m-2)/m}=f_{F}\,.$ (4.31)
@@
$h^{2}\Omega_{\rm{GW}}^{\rm{(1,peak)}}\simeq A_{\Delta}\left(\frac{t_{kin}}{t_{F}}\right)^{2(1+p-4/m)}\,.$ (4.32)
@@
$D=\frac{6(m-2)}{m(2-p)}\,,\qquad E=\frac{4}{n}-\frac{5q}{2}+\frac{4-q\,n}{n(2-p)}\left(\frac{6}{m}+\frac{p}{2}-4\right)\,$, (4.33)
The integral is again dominated by early times (namely E<1E<-1) for kination (regardless of whether it is followed by a matter or radiation period, this being a tracker or not), while for the tracker periods it is dominated by late times. Thus, in the case where only the tracker is present one would need to compare the early and late time contribution to compute the frequency behaviour, etc. We postpone this analysis for future work and focus in the case of volume modulus kination for concreteness, where the earliest times dominate. Let us also point out, however, that in these scenarios the spectral index will be generically different than in scenarios with constant tension. In the case of volume modulus kination, for instance, $\Omega_{\rm{GW}}\sim f^{4}$, resulting in a very large boost in the signal.
@@ 5 Benchmark scenarios
We now turn to illustrating the GW spectra arising from cosmic superstrings in the two scenarios described in Sec. 3.1.2. We operate under the assumption that the PTA signal is explained by cosmic superstrings, and search for further consequences of this assumption [^35] One may instead consider, given the scenarios, what are the values of microscopic parameters that are allowed by current data. We intend to explore this in more generality in future work.. Following Ref. [7], where $\beta=1/2$ and $\chi=0$, the preferred string tension and intercommutation probability are [^36] We note here that in [19] the choice is $\beta=2/3$ and $\chi\neq 0$, which is why they find a different best-fit than in [7].
$G\mu=2\times 10^{-12}\,,\qquad P=4\times 10^{-3}\,.$ (5.1)
Using Eqs. (2.1) and (2.4), this fixes the volume and the string coupling to
$\mathcal{V}=3.3\times 10^{9}\,,\qquad g_{s}=0.11\,$, (5.2)
@@
The vanilla LVS makes a concrete prediction in the LISA band in terms of the mass and decay rate of the volume mode. The transition frequency depends on model-dependent order one numbers, with the most likely scenario being that the LISA mission would be able to measure a deviation from standard cosmology. Due to the large suppression of the signal from a long period of early matter domination, there is no meaningful enhancement in the high-frequency portion of the spectrum (despite an epoch of early kination). The SVL-LVS has an interesting behaviour: the high-frequency signal grows for larger hierarchies between the inflation and late-time vacuum, with a long period of kination in between. As we will discuss, it is not easy to explain the PTA signal whilst being consistent with a large signal at high-frequencies. We nevertheless provide an example where such a large, high-frequency signal can occur as an illustration of the interesting physics behind the fact that the signals grow with large hierarchies. The basic physics can be applied to other well-motivated scenarios which can potentially lead to large signals (see for example [80]).
@@ 5.1 Vanilla LVS and LISA
Consider the mass of the volume modulus in LVS [10]: [^37] Following [158], we define $W_{0}=\sqrt{\frac{2}{\pi}}|e^{K_{\rm{cs}}/2}W_{\rm{GVW}}|$ in terms of the Kähler potential of the complex structure moduli and the Gukov-Vafa-Witten [159] superpotential. This is because [158] finds a $W_{0}$ following a Gaussian distribution around 0.
$m_{\Phi}=\mathcal{O}(1)\times\frac{W_{0}}{\mathcal{V}^{3/2}g_{s}^{1/4}}M_{P}$ (5.3)
which results in a gravitational decay rate, using Eq. (3.20):
$\Gamma_{\Phi}=\frac{\mathcal{O}(1)}{48\pi}\cdot\frac{W_{0}^{3}}{\mathcal{V}^{9/2}g_{s}^{3/4}}M_{P}\,.$ (5.4)
This decay rate then translates into a time for decay of the modulus, which in turn predicts a deviation from the flat spectrum at a frequency that we compute as follows. Because, as discussed in Sec. 4.1, the largest contribution to the amplitude at a frequency $f$ from loops that are formed during matter but source during radiation arises from the latest times, the spectrum at $f$ is dominated by the loops that are formed and source during radiation domination. Out of these, the largest contribution arises at $t_{M}(f)$ as defined in Eq. (4.14). The transition from the flat spectrum sourced by loops that are created and emit during radiation domination is therefore predicted to occur when $t_{i}(t_{M}(f),f)>1/\Gamma_{\Phi}$, meaning that no loop created during radiation domination sources at its point of maximum emission $t_{M}(f)$ in this frequency bin. In formulae, using Eqs. (4.10) and (4.14), this implies a transition $f_{\rm{T}}$ given by:
$f_{\rm{T}}=f_{\Gamma}\left(\frac{\alpha P^{\chi}}{\Gamma G\mu}\right)^{1/2}=\frac{2}{\left(\alpha P^{\chi}\times\Gamma G\mu\right)^{1/2}}\frac{\Gamma_{\Phi}}{z_{\Gamma}}\,$, (5.5)
where the second equality defines $f_{\Gamma}$ as a particular case of $f_{\Delta}$ (cf: Eq. (4.10)) in terms of the time of reheating $t_{\Gamma}=1/\Gamma_{\Phi}$ and its redshift, $z_{\Gamma}$. Assuming instantaneous reheating and standard cosmology after volume decay:
$f_{\rm{T}}=8.6\times 10^{-4}\text{ Hz }\,\times\,\left(\frac{0.1}{\alpha P^{\chi}}\times\frac{2\times 10^{-12}}{G\mu}\times\frac{c_{i}}{48\pi}\right)^{1/2}\times\left(\frac{m_{\Phi}}{30\text{ TeV}}\right)^{3/2}\,$, (5.6)
@@
Figure 2 illustrates the difference between the spectrum for standard cosmology and in vanilla LVS for the benchmark value $m_{\Phi}=30$ TeV, corresponding in LVS to the values in Eq. (5.2), together with $W_{0}\simeq 1$ and modulo order one numbers. The Figure also illustrates the mass-dependence of the frequency with the mass of the modulus. A lighter volume mode corresponds to a lower-frequency signal (since the decay occurs later in time), although in vanilla scenarios these values might be in tension with achieving a successful Nucleosynthesis, as discussed in Sec. 3.1.2. Ignoring this issue, a lighter volume mode can be achieved by choice of $W_{0}$ or by a stabilization mechanism leading to $m_{\Phi}\sim 1/\mathcal{V}^{a}$ with $a>3/2$. A heavier volume, on the other hand, corresponds to a larger transition frequency. Microscopically, this could be obtained for large values of $W_{0}$ (though note that $W_{0}$ is bounded by above for consistent description of the effective theory, see e.g. [160]), or by a stabilization mechanism similar to the one discussed in Appendix A. In addition, Figure 3 zooms into the interferometer band.
Figure 2: GW spectrum sourced by cosmic superstrings with $G\mu=2\times 10^{-12}$ and $P=4\times 10^{-3}$, in the vanilla LVS scenario, discussed in Section 5.1. A period of early matter domination induced by the volume modulus ends by gravitational decay. This induces a dip in the spectrum that will be robustly probed by the LISA mission. A reasonable estimate of the volume mass is 30 TeV (solid black) and is related to the dip frequency by Eq. (5.6). We also show the spectrum in “Standard” cosmology, i.e: radiation domination with constant tension.
@@
There are a multitude of presently operating and proposed GW experiments that could detect the SGWB studied in this Section [^38] We have used the publicly available data from [161] to plot the experimental sensitivities. (see [162] for a review in sources and detection efforts of cosmological GW backgrounds and [40, 42] for proposals at frequencies above the kHz). In the plots, we show sensitivity curves for the following detectors:
  • Ground based GW laser interferometers: LIGO/VIRGO/KAGRA, including both latest bounds [163] and prospect sensitivities [164, 1, 165], and the Einstein Telescope (ET) [166, 167]. Other proposals include the Cosmic Explorer (CE) [168].
  • Space based GW laser interferometers: LISA [169, 170, 171] (see [172] for prospects on cosmic strings). Other proposals include the Big-Bang Observatory (BBO) [173], the Deci-Hertz Interferometer Gravitational-wave Observatory (DECIGO) [174], Taiji [175], TianQin [176] and $\mu$-ARES [177].
  • Atomic interferometers: AION [178, 179] (ground-based). Other (space-based) proposals include MAGIS [180] and AEDGE [181].
  • Pulsar Timing Arrays (PTA): We use the data from the North American Nanohertz Observatory for Gravitational Waves (NANOGrav) [3] and European Pulsar Timing Array (EPTA) [4] collaborations (the latter combined with data from the Indian Pulsar Timing Array [182]). Other collaborations which have confirmed the detection of the SGWB are the Parkes Pulsar Timing Array (PPTA) [5] and the Chinese Pulsar Timing Array (CPTA) [6]. We also include the sensitivity curves from the Square Kilometre Array (SKA) [183].
Figure 3: GW spectrum sourced by cosmic superstrings with $G\mu=2\times 10^{-12}$ and $P=4\times 10^{-3}$ in the large-scale interferometer band for the vanilla LVS scenario, discussed in Section 5.1. A period of early matter domination induced by the volume modulus ends by gravitational decay. This induces a dip in the spectrum that will be robustly probed by the LISA mission. A reasonable estimate of the volume mass is 30 TeV (solid black) and is related to the dip frequency by Eq. (5.6). We also show the spectrum in “Standard” cosmology, i.e: radiation domination with constant tension.